### v1

@@ 3.11 Masking Strategy and Objective Function
A body of works have contributed to both the masking strategy and the objective function. These works begin by either adopting a novel masking strategy that either excels in hiding the salient information [14, 151, 25], or adapting the masking, given a different input type [152, 153]. Further, the objective function is altered to be more suitable for recovering the original input source.
Figure 15: The pipeline of MST as proposed by [14]. The masking is based on the attention maps provided by the teacher.
The Masked Self-Supervised Transformer (MST) [14] selects patches for masking based on low responses as determined by attention maps from a teacher network, which is an EMA of the student network. These selected patches are replaced with a special token. The training of MST includes a dual-objective approach: a reconstruction loss to rebuild the masked inputs, and a cross-entropy loss designed to synchronize the teacher and student networks, particularly in class differentiation. A detailed overview of the pipeline is depicted in Figure 15.
[152] presented a method inspired by BERT to pre-train transformers from point cloud data. First, the points are partitioned into sub-clouds, and a PointNet is applied to extract point embeddings. With the resulting embeddings, the authors trained a discrete Variational Autoencoder (dVAE), where the encoder is a tokenizer, mapping the continuous embeddings into discrete tokens. After this stage, the transformer pre-training is performed, in which the model receives a masked sequence of point embeddings and learns to output the discrete tokens. The masked tokens are replaced by a learnable token.
Masked Scene Contrast [24] is a framework for 3D representation learning that utilizes contrastive learning. This framework generates input pairs through a series of data augmentation techniques and applies complementary masking. The contrastive learning objective is then employed between the features of the unmasked and masked patches. Additionally, the framework incorporates a reconstruction loss to enhance learning efficacy.
[151] tailored the MIM pre-training strategy for human perception tasks. They began by detecting the human parts and masking the patches corresponding to these parts, the objective being that of reconstructing the masked tokens from the visible ones. Besides this, they also generated another masked view of the input image (sampling other human parts) with the objective of aligning both views global representations (by applying a contrastive loss on the [CLS] tokens).
@@
[153] integrated MAE in their method for self-supervised video hashing. A video clip is firstly downsampled using a CNN, and then, a token is extracted for each frame. Two different subsets of token frames are sampled, each being then fed into an encoder and hashed. After some of the hashed embeddings are masked, a decoder is used to reconstruct the original frame tokens. A contrastive loss is used as well to maximize the similarity of the mean hash embeddings between the two subsets.
@@
[25] introduced some improvements to the student-teacher MIM with contrastive pre-training framework. The first addition is a ranking component that divides the patches into two: the ones that contain salient information and the meaningless ones. The former subset is passed to the student, while the latter is given to the teacher, both being masked and leaving the rest visible. The first objective is to reconstruct the masked patches of the student, as they are harder to predict due to containing more contextual information. A second loss is used to align the globally encoded representations of the two subsets by maximizing the similarity between the embeddings of the [CLS] token. While the same encoder is used for both models, the gradients do not flow through the teacher.
Taking a different masking SSL approach, [154] proposed a method whose objective is to be more robust to the downstream task. On the input image, two different sets of augmentations are applied and each view is encoded. Then, a projection layer transforms the encoded representations into multivariate Gaussian distributions. Finally, a series of learnable masks are applied, and for all, the difference between both masked probabilistic embeddings is computed, the objective being that minimizing this difference.
@@
[19] presented a pre-training framework that combines MIM and contrastive learning. Different subsets (a more aggressive one and a lighter one) of augmentations are applied to the input image to create two views, the latter having a portion of its tokens masked. Then, two vision transformer models (where the one for the unmasked image is an EMA of the other model) are used to encode the tokenized patches. The contrastive loss is applied only between the corresponding corrupted patches.
[155] slightly modified the MAE model by adding Gaussian noise to all pixels in the input image. Given this, besides the reconstruction loss of the masked regions, the objective is further extended to decode all denoised patches. The experiments demonstrated the superiority of this method over the original framework.
@@ 3.12 Masking Strategy and Downstream Task
Some studies that have been pivotal to both the masking strategy and the downstream tasks which they were applied to. These papers start by applying MIM on a new task, but due to the different nature of the studied task, the masking strategy is adapted as well.
@@
LEMaRT [158] is an effective pre-training framework when applied on image harmonization as a downstream task. In this approach, the masked patches are replaced with the patches taken from a perturbed version of the original image. The research also investigated what is the best strategy for creating the masks, concluding that random masking works best for image harmonization.
[160] used masked autoencoders to learn rich, generic, transferable and robust facial representations from face videos. The masking prioritizes specific tokens (those containing eyes, nose, mouth and hair). The learned representations are then tested on downstream tasks, such as facial attribute recognition, facial expression recognition, DeepFake detection, and lip synchronization.
The Saturated Mask AutoEncoder (SMAE), introduced by [161], is a two stage approach for few-shot HDR imaging. The first stage focuses on representation learning, which is performed via masked image modeling. In this stage, the method creates two additional images from the original frame using exposure adjustment. Next, all three images are randomly masked with a high masking ratio before passing them to the model.
Aiming to improve the performance on WSI classification, [162] studied several masking strategies to create a hard mining method useful for multiple instance learning. They observed that the best candidates for masking are the salient patches. To identify this type of patches during training, the authors propose a pipeline in which the attention scores provided by a teacher network serve as indicators of patch saliency.
[163] demonstrated that MAEs can be used in class incremental learning as a rehearsal-based method. The efficiency of MAEs, which require only a few patches, allows for the storage of more examples from previous tasks. In addition, the authors design a two-branch MAE architecture to ensure higher quality reconstructions, one branch being responsible with inserting details in the image.
[88] introduced a novel self-supervised pre-training technique designed for point cloud videos, employing a unique approach that involves masking point tubes. This method focused on training the model to accurately reconstruct these masked tubes. Simultaneously, it engaged in training on a temporal cardinality difference task. This dual training strategy enhances the model’s ability to understand both the spatial structure and temporal dynamics inherent in point cloud videos.
Magnetic Resonance scans are generated in k-space and then transformed in the image domain with the inverse Fourier transform. To obtain an image with high quality and fidelity, the k-space needs to be fully sampled, but this is not realistic in most scenarios. To this end, [164] proposed to leverage the MAE framework by masking data in the k-space, represented in 3D: height, width and time, along the first dimension. Reconstructing the missing tokens via the $l_{1}$ loss, the adopted ViT models learn a rich feature representation that is able to estimate the unsampled data from k-space at inference. The estimated k-space is further refined using three sequential transformer-based decoders (one for each pair of dimensions), and employing the High-Dynamic Range loss after each decoder.
[165] addressed the shortcomings of gallbladder cancer detection in static images, proposing the use of video sequences instead. They adopted MAE as a pretext task, but presented an improved masking strategy that is able to hide the malignant regions more consistently, and thus learn a better representation of the disease. The masking strategy involved a Region Selection Network that generates a probability for each token, which is then used to sample the visible tokens.
@@ 3.13 Target Features and Downstream Task
A number of papers have made significant advancements to the targeted downstream tasks, which also implied a contribution to the target features. Most of these works belong to the medical domain [166, 167, 77], and the target features are chosen to be more representative for the input data. Other papers in this category utilize videos as the input signal [168].
[166] proposed a method to pre-train a model using MIM, which is able to process both 2D and 3D ophthalmic images. The authors developed a new module, called Unified Patch Embedding, consisting of two branches, one for each data type. The module divides the inputs into equal patches and then masks a great portion of them. Then, a common encoder computes the latent representations of the visible patches. Finally, two decoders are employed: one that reconstructs the patches, and another one that estimates their gradient maps (composed of horizontal and vertical edge maps). The experiments showed that the method yields state-of-the-art performance in ophthalmic image classification.
@@
Figure 16: The knowledge distillation pipeline proposed by [168].
[168] introduced Masked Video Distillation (MVD), a new method for self-supervised video representation learning, depicted in Figure 16. This approach has two stages. The first part trains masked image models and masked video models as teachers. The second part trains a student with the representations learned by the teachers as target vectors. The masking is also used in this latter stage.
To boost the performance gains of a MAE in the ultrasound imaging domain, [167] introduced an additional task during the pre-training stage. Due to the high noise-to-signal ratios in such images, they are initially blurred, so that masked patches are reconstructed, while the visible patches are deblurred. Thus, in contrast to the original MAE, all patches are passed through the encoder. The experiments on the downstream task of thyroid ultrasound image classification demonstrated leading results.
The task of trajectory prediction heavily suffers from data distribution shifts. To this end, [170] proposed an online training strategy (i.e. test-time adaptation) based on the MAE framework. The reconstruction objective of the masked input represents just an auxiliary task, in addition to the main regression loss on the previously collected states. The authors conjectured that this method helps deeper layers to be effectively optimized, compared with just using the primary objective function.
Motivated by the scarcity of annotated datasets containing medical scans, [77] introduced a novel unsupervised domain adaptation framework based on MAE. The masked reconstruction task is applied to two input signals that are created from two volumetric scan: a local sub-volume and a global downsampled scan. Different from MAE, the authors employed a convolutional architecture. Finally, when applied to the segmentation downstream task, the method uses a teacher-student framework: the teacher (an EMA of the student) generates a pseudo-label segmentation mask of a target domain image, and then, the student is trained using the resulting pair along with a sample from the source domain.
@@ 3.14 Model Architecture and Target Features
Several papers have been influential in both the model architecture and the target features. The modifications presented by these works are tightly coupled. One the one hand, a target feature is proposed, and in order to accommodate it, architectural changes are adopted [171, 172, 173]. On the other hand, the other works are extending the original frameworks [174, 175], or even introducing new pre-training frameworks based on MIM [176, 177].
[174] proposed a couple of modifications to the original MAE pre-training. They used two decoders: one for image reconstruction (the original task) and one for feature representation estimation. The latter one tries to predict the feature representation of the masked patches, the ground-truth coming from an EMA replica of the MAE encoder. For both decoders, some information about the visible patches from the encoder is injected with the cross-attention layers: the former receives low-level context, while the latter is given high-level features.
[171] used as target values the HOG features of the masked regions. The masked patches are replaced with learnable tokens and the architecture is based on a single encoder followed by a linear head for predicting the HOG features.
@@
Masked Shape Prediction (MSP) [172] used geometric shape as prediction target for masked points. This target includes explicit shape context and deep shape features. Additionally, the architecture of MSP modifies cross-attention and self-attention layers to avoid possible masked shape leakage caused by including the masked part positions in the token embeddings. These modifications are designed to restrict the interaction among masked points.
[175] presented a knowledge distillation technique suitable for object detection, based on the masked autoencoder framework. The student network receives a masked image and learns to recover the missing multi-scale features provided by the teacher network. The student and the teacher networks are based on convolutional layers. Thus, in the design of the student, the authors used masked convolutions to avoid the change of masked patches.
@@
Observing that masked pre-training negatively affects the final layers of a deep ViT model, [173] proposed Masked Image Residual Learning (MIRL). The framework consists of dividing a ViT along the depth into an even number of stages. A decoder is added for each stage, which is fed with the corresponding intermediate embeddings, as well as the masked tokens. While the decoders in the first half reconstruct the main components of the input image, the rest estimate the residuals, i.e. the differences between the target and the prediction.
@@ 3.15 Objective function, Downstream Task and Theoretical Analysis
Two studies had an impact on the objective function and the downstream task, while also conducting a theoretical analysis. These works begin by conducting a deep analysis about different MIM aspects, and then they provide solutions in order to improve them.
[26] unveiled several theoretical insights into the MAE paradigm. Initially, it uncovered a link between the MAE framework and contrastive learning principles, revealing that the reconstruction loss in MAE is analogous to the alignment loss found in contrastive learning. Further exploration provided theoretical assurances regarding the downstream efficacy of MAE models. The connection between contrastive learning also implies the presence of feature collapse, a common challenge in contrastive learning, where aligning solely positive samples diminishes model effectiveness. The researchers introduced Uniformity-enhanced MAE as a solution for the feature collapse problem. This adaptation modifies the loss objective to integrate a novel loss function, specifically designed to reduce feature similarity across unmasked views, thereby preserving feature diversity and enhancing model robustness.

### v2

@@ 3.11 Masking Strategy and Objective Function
A body of works have contributed to both the masking strategy and the objective function. These works either introduced a novel masking strategy that excels in hiding the salient information [1, 165, 33], or adapted the masking strategy to a different input type [105, 166]. Further, to maximize the benefit of using a new masking strategy, the objective function is altered to be more suitable for recovering the original input source [105, 32, 33, 167]. For example, [105] performed the masking in the discrete space of a discrete Variational Autoencoder (dVAE) and, due to this change, the reconstruction loss is the cross-entropy. Another example is the work of [166], in which the contrastive loss is employed to align the token representations of two different masked versions of the same image.
Figure 18: MST pipeline, courtesy of [1] (image licensed under CC BY 4.0). MST leverages an ${L}^{1}$ loss to reconstruct the patches that are masked according to the attention maps provided by the teacher, thus falling into the category of reconstruction-based methods depicted in Figure 2. The cross-entropy loss is only used to synchronize the teacher and the student.
@@
A consistent number of papers in this section employed an objective function composed of multiple losses. Masked Scene Contrast [32] is a framework for 3D representation learning that utilizes contrastive learning. This framework generates input pairs through a series of data augmentation techniques and applies complementary masking. The contrastive learning objective is then employed between the features of the unmasked and masked patches. Additionally, the framework incorporates a reconstruction loss to enhance learning efficacy. [166] integrated MAE in their method for self-supervised video hashing. A video clip is first downsampled using a CNN, and then, a token is extracted for each frame. Two different subsets of token frames are sampled, each being then fed into an encoder and hashed. After some of the hashed embeddings are masked, a decoder is used to reconstruct the original frame tokens. A contrastive loss is added to maximize the similarity of the mean hash embeddings between the two subsets.
@@
[167] slightly modified the MAE model by adding Gaussian noise to all pixels in the input image. With this change, besides reconstructing the masked regions, the objective is further extended to decode all denoised patches. The experiments demonstrated the superiority of this method over the original MAE. [33] introduced some improvements to the teacher-student MIM based on contrastive pre-training. The first addition is a ranking component that divides the patches into two: the ones that contain salient information and the meaningless ones. The former subset is passed to the student, while the latter is given to the teacher, both being masked and leaving the rest visible. The first objective is to reconstruct the masked patches of the student, as they are harder to predict due to containing more salient information. A second loss is used to align the globally encoded representations of the two subsets by maximizing the similarity between the embeddings of the [CLS] token. While the same encoder is used for both models, the gradients do not flow through the teacher.
Aside from multi-component losses, informed masking strategies were also employed in some studies. The Masked Self-Supervised Transformer (MST) [1] selects patches for masking based on low responses as determined by attention maps from a teacher network, which is an EMA of the student network. The selected patches are replaced with a special token. The training of MST includes a dual-objective approach: a reconstruction loss to rebuild the masked inputs, and a cross-entropy loss designed to synchronize the teacher and student networks, particularly in class differentiation. A detailed overview of the pipeline is depicted in Figure 18. [165] tailored the MIM pre-training strategy for human perception tasks. They began by detecting the human parts and masking the patches corresponding to these parts, the objective being that of reconstructing the masked tokens from the visible ones. The authors also generated another masked view of the input image (sampling other human parts), aiming to align the global representations of both views (by applying a contrastive loss on the [CLS] tokens).
[5] introduced a masking strategy that selects important patches according to their similarity with a given text, computed in the CLIP embedding space. They further employ a contrastive loss to align text features with the features extracted from the unmasked patches.
[105] presented a method inspired by BERT to pre-train transformers from point cloud data. First, the points are partitioned into sub-clouds, and a PointNet is applied to extract point embeddings. With the resulting embeddings, the authors trained a dVAE, where the encoder is a tokenizer, mapping the continuous embeddings into discrete tokens. After this stage, the transformer pre-training is performed, in which the model receives a masked sequence of point embeddings and learns to output the discrete tokens. The masked tokens are replaced by a learnable token.
@@
A couple of papers opted for a contrastive objective between two different views of an image obtained by distinct augmentation. Taking a different masking SSL approach, [168] proposed a method whose objective is to be more robust for downstream tasks. On the input image, two different sets of augmentations are applied. The augmented images are encoded, and a projection layer transforms the encoded representations into multivariate Gaussian distributions. Finally, a series of learnable masks are applied. The difference between both masked probabilistic embeddings is computed, the objective being that of minimizing this difference. [3] presented a pre-training framework that combines MIM and contrastive learning. Different subsets (a more aggressive one and a lighter one) of augmentations are applied to the input image to create two views, the latter having a portion of its tokens masked. Then, two vision transformer models (where the one for the unmasked image is an EMA of the other model) are used to encode the tokenized patches. The contrastive loss is applied only between the corresponding corrupted patches.
@@ 3.12 Masking Strategy and Downstream Task
The research highlighted in this subsection focuses on developing MIM pre-training methods that are effective on specific downstream tasks. These studies accomplish this by creating custom masking strategies [169, 170, 106, 171] or by transforming the input signal into a specialized feature space [172, 173, 174, 175, 176]. The custom masking strategies are created such that the features of interest are masked in the pre-training phase. For example, [169] learn representations that are meaningful for facial recognition by masking the regions that contain the eyes, nose and mouth, during pre-training.
To improve the masking strategy, some papers proposed an informed masking policy. [169] used masked autoencoders to learn rich, generic, transferable and robust facial representations from face videos. The masking prioritizes specific tokens (those containing eyes, nose, mouth and hair). The learned representations are then tested on downstream tasks, such as facial attribute recognition, facial expression recognition, DeepFake detection, and lip synchronization. Aiming to improve the performance on WSI classification, [177] studied several masking strategies to create a hard mining method useful for multiple instance learning. They observed that the best candidates for masking are the salient patches. To identify this type of patches during training, the authors proposed a pipeline in which the attention scores provided by a teacher network serve as indicators of patch saliency. [171] addressed the shortcomings of gallbladder cancer detection in static images, proposing the use of video sequences instead. They adopted MAE as a pretext task, but presented an improved masking strategy that is able to hide the malignant regions more consistently, and thus learn a better representation of the disease. The masking strategy involved a Region Selection Network that generates a probability for each token, which is then used to sample the visible tokens.
A few papers adopted MIM for increasing the quality of images. The Saturated Mask AutoEncoder (SMAE), introduced by [175], is a two-stage approach for few-shot HDR imaging. The first stage focuses on representation learning, which is performed via masked image modeling. In this stage, the method creates two additional images from the original frame using exposure adjustment. Next, all three images are randomly masked with a high masking ratio before passing them to the model. LEMaRT [173] is an effective pre-training framework when applied on image harmonization as a downstream task. In this approach, the masked patches are replaced with the patches taken from a perturbed version of the original image. The research also investigated what is the best strategy for creating the masks, concluding that random masking works best for image harmonization. Magnetic Resonance scans are generated in k-space and then transformed in the image domain with the inverse Fourier transform. To obtain an image with high quality and fidelity, the k-space needs to be fully sampled, but this is not realistic in most scenarios. To this end, [176] proposed to leverage the MAE framework by masking data in the k-space, represented in 3D: height, width and time, along the first dimension. By reconstructing the missing tokens via the $L^{1}$ loss, the adopted ViT models learn a rich feature representation that is able to estimate the unsampled data from k-space at inference. The estimated k-space is further refined using three sequential transformer-based decoders (one for each pair of dimensions), and employing the High-Dynamic Range loss after each decoder.
[178] demonstrated that MAEs can be used in class incremental learning as a rehearsal-based method. The efficiency of MAEs, which require only a few patches, allows for the storage of more examples from previous tasks. In addition, the authors designed a two-branch MAE architecture to ensure higher quality reconstructions, one branch being responsible with inserting details into the image.
@@
[106] introduced a novel self-supervised pre-training technique designed for point cloud videos, employing a unique approach that involves masking point tubes. This method focused on training the model to accurately reconstruct these masked tubes. Simultaneously, the model is engaged in training on a temporal cardinality difference task. This dual training strategy enhances the model’s ability to understand both the spatial structure and temporal dynamics inherent in point cloud videos.
@@ 3.13 Target Features and Downstream Task
Several studies aiming to develop MIM methods optimized for specific downstream tasks achieved their goal by exploring various types of target features for masking. Most of these works are in the medical domain [179, 180, 181]. The remaining two are unrelated, MAGVLT [182] being a visual-language transformer, and MVD [183] being a distillation framework.
The adoption of MIM in the medical domain is highly motivated by the scarcity of annotated datasets. [179] proposed a method to pre-train a model using MIM, which is able to process both 2D and 3D ophthalmic images. The authors developed a new module, called Unified Patch Embedding, consisting of two branches, one for each data type. The module divides the inputs into equal patches and then masks most of them. Then, a common encoder computes the latent representations of the visible patches. Finally, two decoders are employed: one that reconstructs the patches, and another one that estimates their gradient maps (composed of horizontal and vertical edge maps). The experiments showed that the method yields state-of-the-art performance in ophthalmic image classification. To boost the performance gains of an MAE in the ultrasound imaging domain, [180] introduced an additional task during the pre-training stage. Due to the high noise-to-signal ratios in such images, they are initially blurred, so that masked patches are reconstructed, while the visible patches are deblurred. Thus, in contrast to the original MAE, all patches are passed through the encoder. The experiments on the downstream task of thyroid ultrasound image classification demonstrated leading results. Different from the previous methods, [181] introduced a novel unsupervised domain adaptation framework based on MAE, in which the authors employed a convolutional architecture. The masked reconstruction task is applied to two input signals that are created from two volumetric scans: a local sub-volume and a global downsampled scan. When applied to the segmentation downstream task, the method uses a teacher-student framework. The teacher (an EMA of the student) generates a pseudo-label segmentation mask of a target domain image, and then, the student is trained using the resulting pair along with a sample from the source domain.
@@
Figure 19: The MVD knowledge distillation pipeline, courtesy of [183] (image available under the ArXiv perpetual license agreement). MVD is classified as a reconstruction-based approach (belongs to the generic pipeline illustrated in Figure 2), as it uses an $L^{2}$ loss between the output features of its two decoders and the features provided by two teacher networks, one for video and one for images.
[183] introduced Masked Video Distillation (MVD), a new method for self-supervised video representation learning, depicted in Figure 19. This approach has two stages. The first stage is to train masked image models and masked video models as teachers. The second stage is to train a student with the representations learned by the teachers as target vectors. The masking is also used in this latter stage.
@@ 3.14 Model Architecture and Target Features
Some studies contributed in terms of both model architecture and target features. The modifications proposed by these works are tightly coupled. On the one hand, some works proposed new target features and consequently implemented architectural adjustments to support these features [184, 185]. On the other hand, other works extended the existing MIM frameworks [186, 187], or even introduced new pre-training frameworks based on MIM [188, 189].
Rather than using a single decoder, some works introduced multiple decoders in their pipeline. [186] proposed a couple of modifications to the original MAE pre-training. They used two decoders: one for image reconstruction (the original task) and one for feature representation estimation. The latter one tries to predict the feature representation of the masked patches, the ground-truth coming from an EMA replica of the MAE encoder. For both decoders, some information about the visible patches from the encoder is injected with the cross-attention layers: the former receives low-level context, while the latter is given high-level features. Observing that masked pre-training negatively affects the final layers of a deep ViT model, [185] proposed Masked Image Residual Learning (MIRL). The framework consists of dividing a ViT along the depth into an even number of stages. A decoder is added for each stage, which is fed with the corresponding intermediate embeddings, as well as the masked tokens. While the decoders in the first half reconstruct the main components of the input image, the rest estimate the residuals, i.e. the differences between the target and the prediction.
@@
Masked Shape Prediction (MSP) [184] uses geometric shape as prediction target for masked points. This target includes explicit shape context and deep shape features. Additionally, the architecture of MSP modifies cross-attention and self-attention layers to avoid possible masked shape leakage caused by including the masked part positions in the token embeddings. These modifications are designed to restrict the interaction among masked points.
[187] presented a knowledge distillation technique suitable for object detection, based on the MAE framework. The student network receives a masked image and learns to recover the missing multi-scale features provided by the teacher network. The student and the teacher networks are based on convolutional layers. Thus, in the design of the student, the authors used masked convolutions to avoid information leakage.
@@
[190] proposed to integrate masked modeling for pre-training video-language tasks. A common cross-modal transformer is employed for the tokens of both modalities, with the objective of reconstructing them. While the text tokens follow the approach in BERT, the video frames are tokenized into discrete values with a dVAE, the masking being achieved at the pixel level by replacing image patches with zero. Their experiments attest the benefits of reconstructing discrete visual tokens.
@@ 3.15 Objective function, Downstream Task and Theoretical Analysis
Two studies had an impact on the objective function and the downstream task, while also conducting a theoretical analysis. These works [34, 191] begin by conducting a deep analysis about different MIM aspects, and then, they provide solutions in order to improve them [34] or exploit the properties induced by MIM to implement a particular downstream task [191].
[34] unveiled several theoretical insights of the MAE paradigm. Initially, they uncovered a link between the MAE framework and contrastive learning principles, revealing that the reconstruction loss in MAE is analogous to the alignment loss found in contrastive learning. Further exploration provided theoretical assurances regarding the downstream efficacy of MAE models. The connection with contrastive learning also implies the presence of feature collapse, a common challenge in contrastive learning, where aligning solely positive samples diminishes model effectiveness. The researchers introduced Uniformity-enhanced MAE as a solution for the feature collapse problem. This adaptation modifies the loss objective to integrate a novel loss function, specifically designed to reduce feature similarity across unmasked views, thereby preserving feature diversity and enhancing model robustness.